You can determine the index of refraction of a substance by measuring its critical angle for total internal reflection.

Randomized Variables
Θc,w = 74.6°

(a) What is the index of refraction of a substance that has a critical angle of 74.6° when submerged in water (with index of refraction 1.333)?
(b) What would the critical angle be for this substance in air?
1
Expert's answer
2020-11-20T07:13:17-0500

(a)

"n_1 = ?"

"n_2 = 1.333"

"\u03b8_c = 74.6\u00ba"

"n_1 = \frac{n_2}{sin(θ_c)}"

"n_1 = \frac{1.333}{sin(74.6º)} = \frac{1.333}{0.9641} = 1.3826"

(b) "θ_c = sin^{-1}\frac{n_2}{n_1}"

"θ_c = sin^{-1}\frac{1.000293}{1.3826}= sin^{-1}(0.72348) = 46.34º"